40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 17 16:50:53 2022, Paco, Update, BHD, XARM AUX alignment 
    Reply  Fri Mar 18 17:15:19 2022, Paco, Update, BHD, XARM AUX alignment cont. Screen_Shot_2022-03-18_at_4.09.38_PM.png
       Reply  Mon Mar 21 18:00:05 2022, Ian MacMillan, Update, BHD, XARM AUX alignment cont. PXL_20220322_015247138.jpg
Message ID: 16735     Entry time: Fri Mar 18 17:15:19 2022     In reply to: 16732     Reply to this: 16739
Author: Paco 
Type: Update 
Category: BHD 
Subject: XARM AUX alignment cont. 

[Paco, Ian]

First, we re-balanced the ITMX OSEMs so that they would damp at around half-a-shadow.


Then, we set up a clean camera inside ITMX chamber looking at the ITMX optic. Then, using the live feed we aligned the AUX beam from the ETMX station using M1 and M2. The camera was great to help us align the beam properly close to the ITMX center. It wasn't very long until we could see a green beam on the IR card, but we didn't really see any flashing, so this may just be the bare transmission away from XARM resonance (Attachment #1).

Ian checked the reflection from ITMX using the IR card with holes, and he pretty much only saw one beam spot, so we turned to look for a beam scattering on the vacuum tube but didn't really see anything. This could mean that we were hitting the ETMX again, or missing slightly, or missing completely. We tried scanning the ITMX pitch and yaw using the bias (alignment) sliders, and with the illuminators off, try seeing some scattered green beam on the ETMX. We can't really see anything yet, but we will keep trying. If there are any tips on our method, it would be great to know them.

Attachment 1: Screen_Shot_2022-03-18_at_4.09.38_PM.png  3.428 MB  Uploaded Fri Mar 18 18:24:02 2022  | Hide | Hide all
Screen_Shot_2022-03-18_at_4.09.38_PM.png
ELOG V3.1.3-